It's algebra
M over 15 + 62 = 23 What is m?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

M over 15 + 62 = 23 What is m?
-------
If M = m:
m/(15+62) = 23
m/77 = 23
m = 23*77
m = 1771

Answer by stanbon(75887)

M over 15 + 62 = 23 What is m?
(m/15) + 62 = 23
---
(m/15) = -39
m = 15*-39
m = -585
==================
Cheers,
Stan H.